# Feynman Graph Integrals on Kähler Manifolds

### v2

@@ Abstract
In this paper, we provide a rigorous proof of the convergence, in the Cauchy principal value sense, of Feynman graph integrals of holomorphic quantum field theories on closed real analytic Kähler manifolds. In the special case of Calabi–Yau manifolds, these graph integrals give rise to new invariants of Calabi–Yau metrics.
@@ 1 Introduction
Feynman graph integrals, originally introduced in physics, have become central to numerous areas of mathematics. A notable example are the configuration space integrals for topological quantum field theories developed by S. Axelrod and I. Singer [2, 3] and by M. Kontsevich [21]. These integrals have proved exceptionally powerful, forming the foundation of modern approaches to knot invariants [21], 3d manifold invariants [7, 6], smooth structures on fiber bundles [21], operad theory [16, 23], deformation quantization of Poisson manifolds [24], and many other applications.
In contrast, a systematic theory of Feynman graph integrals in holomorphic quantum field theories has only begun to take shape in the past decade [12, 25, 27, 28, 26, 8, 36, 34, 33]. A key technical obstacle is that these integrals are typically not absolutely convergent (see [33, Example 2.1.1]). To address this, the first author recently introduced a heat-kernel regularization and established the finiteness of holomorphic field theories on affine spaces; see [33]. See also [11] for further discussion of heat-kernel regularization in renormalization in quantum field theory.
This paper presents an alternative approach to [33], offering another formulation of Feynman graph integrals in holomorphic quantum field theories using Cauchy principal values (see [18]). Our construction extends the setting of [33] from affine spaces to arbitrary closed real analytic Kähler manifolds, thereby broadening the scope of holomorphic quantum field theories.
@@ 1.1 Main results
For simplicity, and to avoid introducing too many concepts and notations at the beginning of the paper, we state here a special case of our main results (Theorem 2.20 and Theorem 2.25).
Let $M$ be a closed Calabi-Yau manifold, whose holomorphic volume form is denoted by $\Omega$.
Let $P$ be a distribution-valued section of the bundle $\Lambda^{\bullet}T_{0,1}^{*}(M\times M)$ over $M\times M$, such that the following equations hold (see Definition 2.5 for details):
@@
where $\delta$ and $\mathcal{H}$ are the integral kernels of the identity operator and the harmonic projection operator, respectively.
@@
$\lim_{\epsilon\rightarrow 0}\int_{M^{\vec{\Gamma}_{0}}}\chi_{\epsilon}(p_{1},\cdots,p_{|\vec{\Gamma}_{0}|})\Big{(}\prod_{e\in\vec{\Gamma}_{1}}P\big{(}p_{h(e)},p_{t(e)}\big{)}\Big{)}\Big{(}\prod_{i\in\vec{\Gamma}_{0}}\Omega(p_{i})\Big{)}$
@@
$\chi_{\epsilon}(p_{1},\cdots,p_{|\vec{\Gamma}_{0}|})=\begin{cases}0,&\text{if }\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}(p_{i},p_{j})\leq\epsilon,\\ 1,&\text{if }\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}(p_{i},p_{j})>\epsilon.\end{cases}$
@@
By the uniqueness of $P$ (see 2.6), it follows that the integrals defined in 1.1 are invariants of the Calabi–Yau metric and the holomorphic volume form.
Using Cauchy principal values in Appendix B and the Fulton–MacPherson compactification of the configuration spaces, we can reformulate Theorem 1.1 as follows:
@@
$\lim_{\epsilon\rightarrow 0}\int_{M^{\vec{\Gamma}_{0}}}\chi_{\epsilon}(p_{1},\dots,p_{|\vec{\Gamma}_{0}|})\Big{(}\prod_{e\in\vec{\Gamma}_{1}}P(p_{h(e)},p_{t(e)})\Big{)}\Big{(}\prod_{i\in\vec{\Gamma}_{0}}\Omega(p_{i})\Big{)}$
@@
$=\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{\widetilde{Conf}_{\vec{\Gamma}_{0}}(M)}\Big{(}\prod_{e\in\vec{\Gamma}_{1}}P(p_{h(e)},p_{t(e)})\Big{)}\Big{(}\prod_{i\in\vec{\Gamma}_{0}}\Omega(p_{i})\Big{)},$
@@
where $\widetilde{Conf}_{\vec{\Gamma}_{0}}(M)$ is the Fulton–MacPherson compactification of the configuration space of $|\vec{\Gamma}_{0}|$ points in $M$ (see 2.14 ), and the integral on the right hand side is in the sense of Cauchy principal values (see Appendix B).
@@
Theorem 1.1 and Theorem 1.2 are special cases of our main results Theorem 2.20 and Theorem 2.25, where we consider general closed real analytic Kähler manifolds and Hermitian holomorphic vector bundles with holomorphic pairings.
@@
Notably, the proof of our main theorems draws on Getzler’s rescaling technique (see [4, 15]), originally developed for the Atiyah–Singer local index theorem, highlighting a fruitful interplay between index-theoretic methods and Feynman graph integrals. In future work, we plan to adapt the machinery employed in the proof of the family version of the Atiyah–Singer local index theorem to the study of Feynman graph integrals.
1.2 Applications and outlook
  1. (1)
    Constructions of geometric invariants as Theorem 1.1.
  2. (2)
    Construction of factorization algebras, as in [34]; see also [9, 10] for foundational discussions on factorization algebras.
  3. (3)
    Constructions of higher chiral algebras as in [17].
  1. (1)
    Gauge field theories on $\mathbb{R}^{4}$. In the well-known Penrose–Ward correspondence (see [31, 35]), it is conjectured that perturbative self-dual Yang-Mills theory should correspond to a perturbative holomorphic quantum field theory on the twistor space. We expect that our theorems offer new insights toward understanding this correspondence. For example, our results should imply the finiteness of Feynman graph integrals in self-dual Yang-Mills theory.
  2. (2)
    Holomorphic “knot” invariants. Feynman graph integrals arising from topological field theories have proven useful in the study of knot invariants. Our theorems have potential applications in the study of embeddings of Riemann surfaces—viewed as holomorphic “knots”—into complex manifolds. For the study of holomorphic linking numbers in the literature, see [1, 13].
  3. (3)
    Mirror Symmetry. In mirror symmetry, the Gromov-Witten invariants of Calabi-Yau manifolds are conjecturally equal to the Feynman graph integrals of Kodaira-Spencer gravity theory (or BCOV theory, see [5]) on the mirror manifolds. Since Kodaira-Spencer gravity theory is a holomorphic quantum field theory, it is natural to expect that our theorems can be applied in this context.
@@ 1.3 Organizations
In Section 2, we introduce the notion of the propagator. Assuming 2.11—that is, that the propagator has the desired singularities, namely (logarithmic) divisorial type singularities (B.12)—we define Feynman graph integrals using Cauchy principal values. This construction forms the core of our main results, Theorem 2.20 and Theorem 2.25.
Section 3 and Section 4 are devoted to the proof of 2.11. In Section 3, we introduce the notion of regular expressions, which can be viewed as the analog of divisorial type singularities in the context of heat kernels. In Section 3.2, we establish the connection between regular expressions of the heat kernel and the divisorial type singularities of the propagator, as stated in 3.14. This reduces the proof of 2.11 to verifying that the heat kernel admits a regular expression, namely 4.3.
@@ 2.1 Propagators
Therefore, $\mathcal{H}^{\prime\prime}$ is a harmonic section of $\widetilde{E}\boxtimes\widetilde{E}^{\prime}$. Since the Laplacian operator is an elliptic operator, we conclude that $\mathcal{H}^{\prime\prime}$ is a smooth section of $\widetilde{E}\boxtimes\widetilde{E}^{\prime}$ by the regularity of elliptic differential equations (see [19]). ∎
@@
The existence and uniqueness of the heat kernel follow from [4]. The convergence $\lim_{t\to\infty}H_{t}=\mathcal{H}$ could be proved as in [4, Section 2.6].
@@
The proof is similar to that of [4, Theorem 2.38], and we outline the main steps.
@@
Since the singularities of the right-hand side are supported on the diagonal $\triangle$, it follows from elliptic regularity (cf. [19]) that $\lim_{\epsilon\to 0}G_{\epsilon,L}$ is smooth off $\triangle$. Hence, so is
@@
By [4, Proposition 2.37], the integral
@@
This follows directly from [4, Proposition 2.37]. ∎
@@
Hence, by 2.9, the singularities of $P$ are fully determined by the singularities of $P_{t}$ at $t=0$.
@@ 2.2 Feynman graph integrals
2.11 is the most technically involved result in this paper. The key ingredients for its proof are presented in section 3 and section 4, while the complete proof is given in section 4.2.1. In this section, we assume the validity of 2.11 and use it to construct Feynman graph integrals.
@@
The analyticity assumption in 2.11 is essential for our proof. It remains an interesting question whether this theorem holds without it.
@@
For a coordinate-free and equivalent definition of holomorphic Lagrangian densities, see [36, Definition 2.15].
@@
Let $(M,E,\omega)$ be a triple as in 2.11. Let $\vec{\Gamma}$ be a directed graph. Then
@@
Let $(M,E,\omega)$ be triples as in 2.11. Let $\vec{\Gamma}$ be a directed graph. The Feynman graph integral on $M$ is defined by:
@@
When $M$ is an elliptic curve and $E$ is the trivial vector bundle, the Feynman graph integrals can be computed explicitly, see [37].
In the remainder of this subsection, we introduce an alternative formulation of 2.20.
@@
Let $(M,E,\omega)$ be triples as in 2.11. Let $\vec{\Gamma}$ be a directed graph, and $\tilde{\rho}$ a fake distance function. Then we have

### v3

@@ Abstract
In this paper, we establish the convergence of Feynman graph integrals on closed real-analytic Kähler manifolds and uncover the structural mechanism underlying this convergence. The key insight is that, using Getzler’s rescaling technique, the graph integrands extend canonically to the Fulton–MacPherson compactification of configuration spaces as forms with divisorial-type singularities. This allows the Feynman graph integrals to be rigorously defined as Cauchy principal value integrals. As an application, these integrals provide a mathematically rigorous construction of the higher-genus B-model invariants on Calabi–Yau threefolds in the sense of Bershadsky–Cecotti–Ooguri–Vafa (BCOV).
@@ 1 Introduction
Feynman graph integrals lie at the core of perturbative quantum field theory and have driven developments across geometry, topology, algebra, and analysis in mathematics. While their algebraic and combinatorial structures are well understood, providing a mathematically rigorous meaning remains subtle. Understanding the convergence of Feynman integrals in certain quantum field theories has far-reaching applications across mathematics and physics. Notable examples include the configuration space integrals for topological quantum field theories developed by S. Axelrod and I. Singer [axelrod1994chern, Axelrod1991ChernSimonsPT] and by M. Kontsevich [kontsevich1994feynman], which have proved exceptionally powerful. These integrals offer modern approaches to knot invariants [kontsevich1994feynman], 3-manifold invariants [bott2018integral, bott1999integral], smooth structures on fiber bundles [kontsevich1994feynman], operad theory [getzler1994operads, kontsevich1999operads], deformation quantization of Poisson manifolds [kontsevich2003deformation], and numerous other applications.
In contrast, a systematic theory of Feynman graph integrals in holomorphic quantum field theories has emerged only in the past decade [Costello2015QuantizationOO, Li:2011mi, li2021regularized, li2023regularized, Li:2016gcb, budzik2023feynman, Williams:2018ows, wang2024factorization, wang2025feynman]. A major technical challenge is that these integrals are typically not absolutely convergent (see [wang2025feynman, Example 2.1.1]). To overcome this, the first author recently introduced a heat-kernel regularization and established the finiteness of holomorphic field theories on affine spaces [wang2025feynman]. For a comprehensive account of heat-kernel regularization , we refer the reader to [costellorenormalization].
This work offers an alternative approach to [wang2025feynman]. We show that the Feynman graph integrand extends canonically to the Fulton–MacPherson compactification of configuration spaces with mild singularities, allowing the corresponding graph integrals to be defined rigorously as Cauchy principal value integrals (c.f. [herrera1971residues]). The resulting convergence theorem extends from affine spaces to arbitrary closed real-analytic Kähler manifolds, thereby broadening the scope of holomorphic quantum field theories.
  1. (1)
    Constructions of invariants as in 1.5.
  2. (2)
    Construction of factorization algebras, as in [wang2024factorization]; see also [costello_gwilliam_2016, costello_gwilliam_2021] for foundational discussions on factorization algebras.
  3. (3)
    Constructions of higher chiral algebras as in [minghaoBrianZhengping2025].
  1. (1)
    Mirror Symmetry. In mirror symmetry, the Gromov–Witten invariants of Calabi–Yau manifolds are conjecturally identified with the Feynman graph integrals arising from the Kodaira–Spencer gravity theory (also known as the BCOV theory, see [bershadsky1994kodaira, costello2012quantum]) on the mirror manifolds. Our results define the higher-genus B-model invariants and provide a rigorous mathematical formulation of the BCOV mirror symmetry conjecture (see also section 1.2).
  2. (2)
    Gauge field theories on $\mathbb{R}^{4}$. In the well-known Penrose–Ward correspondence (see [Penrose:1976js, WARD197781]), it is conjectured that perturbative self-dual Yang-Mills theory should correspond to a perturbative holomorphic quantum field theory on the twistor space. We expect that our theorems offer new insights toward understanding this correspondence. For example, our results should imply the finiteness of Feynman graph integrals in self-dual Yang-Mills theory.
  3. (3)
    Holomorphic “knot” invariants. Feynman graph integrals arising from topological field theories have proven useful in the study of knot invariants. Our theorems have potential applications in the study of embeddings of Riemann surfaces, viewed as holomorphic “knots”, into complex manifolds. For the study of holomorphic linking numbers in the literature, see [Atiyah:1981ey, Frenkel:2005qk].
@@ 1.1 Main results
For simplicity, and to avoid introducing excessive technical notions at the outset, we present here a special case of our main results (Theorems 2.21 and 2.26), which provides a rigorous mathematical construction of the higher-genus B-model in the framework of Bershadsky–Cecotti–Ooguri–Vafa (BCOV).
Let $M$ be a closed Calabi–Yau threefold equipped with a Calabi–Yau metric and a holomorphic volume form $\Omega$. Consider the bundle
$\widetilde{E}=\wedge^{\ast}T^{1,0}M\otimes\wedge^{\ast}(T^{0,1}M)^{\ast},$
whose space of sections is
$\Gamma(\widetilde{E})=\Omega^{0,\ast}\big(M,\wedge^{\ast}T^{1,0}M\big).$
Contraction with $\Omega$ gives an isomorphism
$(-)^{\vee}_{\Omega}:\Omega^{0,j}(M,\wedge^{i}T^{1,0}M)\;\longrightarrow\;\Omega^{3-i,j}(M).$
We view the operators $\partial$ and $\bar{\partial}$ on $\Omega^{\ast,\ast}(M)$ as operators on $\Omega^{0,\ast}(M,\wedge^{\bullet}T^{1,0}M)$ via this identification.
Consider the trace map
$\operatorname{tr}:\Omega^{0,j}(M,\wedge^{i}T^{1,0}M)\to\Omega^{6-i,j}(M),\quad\alpha\mapsto(\alpha^{\vee}_{\Omega})\wedge\Omega.$ (1.1)
Let $P$ be a distribution-valued section of the bundle ${\widetilde{E}}\boxtimes{\widetilde{E}}$ over $M\times M$, such that the following equations hold (see Definition 2.5 and 2.6 for details):
@@
where $\delta$ and $\mathcal{H}$ are the delta-distribution and the integral kernel of the harmonic projection operator, respectively.
The BCOV propagator is given by
$P_{\mathrm{BCOV}}:=(\partial\otimes\mathrm{id})\,P.$
@@
$\lim_{\epsilon\rightarrow 0}\int_{M^{\vec{\Gamma}_{0}}}\chi_{\epsilon}(p_{1},\cdots,p_{|\vec{\Gamma}_{0}|})\left(\bigotimes_{v\in\vec{\Gamma}_{0}}\mathrm{tr}_{v}\right)\left(\prod_{e\in\vec{\Gamma}_{1}}P_{\mathrm{BCOV}}(p_{t(e)},p_{h(e)})\right)$ (1.2)
@@
$\chi_{\epsilon}(p_{1},\cdots,p_{|\vec{\Gamma}_{0}|})=\begin{cases}0,&\text{if }\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}(p_{i},p_{j})\leq\epsilon,\\ 1,&\text{if }\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}(p_{i},p_{j})>\epsilon,\end{cases}$
and for a finite set $A$, $|A|$ denotes number of elements in $A$.
@@
The convergence of the limits in (1.2) arises from a structural property of the Feynman graph integrand, described in 1.2 below. Let $\widetilde{Conf}_{\vec{\Gamma}_{0}}(M)$ denote the Fulton–MacPherson compactification (see 2.15) of the configuration space of $|\vec{\Gamma}_{0}|$ points in $M$. We prove the following:
Theorem 1.2.
Under the assmuption of 1.1, the Feynman graph integrand
$\left(\bigotimes_{v\in\vec{\Gamma}_{0}}\mathrm{tr}_{v}\right)\left(\prod_{e\in\vec{\Gamma}_{1}}P_{\mathrm{BCOV}}(p_{t(e)},p_{h(e)})\right)$
extends canonically to a form on $\widetilde{Conf}_{\vec{\Gamma}_{0}}(M)$ with divisorial-type singularities (B.12) .
Using Cauchy principal value integrals in Appendix B and 1.2, we can reformulate Theorem 1.1 as follows:
@@
$\lim_{\epsilon\rightarrow 0}\int_{M^{\vec{\Gamma}_{0}}}\chi_{\epsilon}(p_{1},\dots,p_{|\vec{\Gamma}_{0}|})\left(\bigotimes_{v\in\vec{\Gamma}_{0}}\mathrm{tr}_{v}\right)\left(\prod_{e\in\vec{\Gamma}_{1}}P_{\mathrm{BCOV}}(p_{t(e)},p_{h(e)})\right)$
@@
$=\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{\widetilde{Conf}_{\vec{\Gamma}_{0}}(M)}\left(\bigotimes_{v\in\vec{\Gamma}_{0}}\mathrm{tr}_{v}\right)\left(\prod_{e\in\vec{\Gamma}_{1}}P_{\mathrm{BCOV}}(p_{t(e)},p_{h(e)})\right),$
@@
where the integral on the right hand side is in the sense of Cauchy principal value integral (see Appendix B).
@@
1.1, 1.2 and 1.3 are special cases of our main results Theorem 2.21 and 2.26, where we consider general closed real analytic Kähler manifolds and Hermitian holomorphic vector bundles with holomorphic pairings.
@@
Notably, the proof of our main theorems draws on Getzler’s rescaling technique (see [berline2003heat, getzler1986short]), originally developed for the Atiyah–Singer local index theorem, highlighting a fruitful interplay between index-theoretic methods and Feynman graph integrals. In future work, we plan to adapt the machinery employed in the proof of the family version of the Atiyah–Singer local index theorem to the study of Feynman graph integrals.
1.2 BCOV’s formulation of Mirror symmetry conjecture
In this section, under the same setting as in section 1.1, we apply our convergence theorems to briefly describe the higher-genus B-model invariants in the BCOV framework and to formulate their mirror symmetry conjecture.
It follows from 1.3 and the graded symmetric property of $P_{\mathrm{BCOV}}$ that
Corollary 1.4.
Let $\Gamma$ be a trivalent undirected graph. Then the integral
$W_{\mathrm{BCOV}}(\Gamma)=\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{\widetilde{\mathrm{Conf}}_{\Gamma}(M)}\left(\bigotimes_{v\in\Gamma_{0}}\mathrm{tr}_{v}\right)\left(\prod_{e\in\Gamma_{1}}P_{\mathrm{BCOV}}(p_{t(e)},p_{h(e)})\right)$
is well-defined in the sense of Cauchy principal value.
Definition 1.5.
Let $\mathrm{Graph}_{3}$ denote the set of connected trivalent undirected graphs. For $g\geq 2$, define
$F_{g}^{\mathrm{BCOV}}(M)=\sum_{\begin{subarray}{c}\Gamma\in\mathrm{Graph}_{3}\\ g(\Gamma)=g\end{subarray}}\frac{1}{|\mathrm{Aut}(\Gamma)|}\,W_{\mathrm{BCOV}}(\Gamma),\qquad g(\Gamma)=|\Gamma_{1}|-|\Gamma_{0}|+1.$
Now we can formulate BCOV conjecture mathematical rigorously as follows:
Conjecture 1.6 (Bershadsky–Cecotti–Ooguri–Vafa).
For $g\geq 2$, $F_{g}^{\mathrm{BCOV}}(M)$ defines the genus-$g$ B-model invariant, i.e.
  1. 1.
    $F_{g}^{\mathrm{BCOV}}(M)$ is independent of the Kähler class of $M$;
  2. 2.
    if $M$ and $M^{\vee}$ are mirror manifolds, then the holomorphic limit (see [kanazawa2015lectures]) of $F_{g}^{\mathrm{BCOV}}(M)$ coincides with the genus-$g$ Gromov–Witten invariants of the mirror $M^{\vee}$.
Remark 1.7.
For the elliptic curve $E_{\tau}$, S. Li and K. Costello [li2011bcov, costello2012quantum] constructed a variant of $F_{g}^{\mathrm{BCOV}}(E_{\tau})$ with punctures, whose large complex structure limit as $\bar{\tau}\to\infty$ (with $\tau$ fixed) recovers the punctured Gromov–Witten invariants.
@@ 1.3 Organizations
In Section 2, we introduce the notion of the propagator. Assuming 2.12, that is, that the propagator has the desired singularities, we define Feynman graph integrals using Cauchy principal value integrals. This construction forms the core of our main results, Theorem 2.21 and Theorem 2.26.
Sections Section 3 and Section 4 are devoted to the proof of 2.12. In Section 3, we introduce the notion of regular expressions, which can be viewed as the analog of divisorial type singularities in the context of heat kernels. In Section 3.2, we establish the connection between regular expressions of the heat kernel and the divisorial type singularities of the propagator, as stated in 3.14. This reduces the proof of 2.12 to verifying that the heat kernel admits a regular expression, namely 4.3.
@@ 2.1 Propagators
Therefore, $\mathcal{H}^{\prime\prime}$ is a harmonic section of $\widetilde{E}\boxtimes\widetilde{E}^{\prime}$. Since the Laplacian operator is an elliptic operator, we conclude that $\mathcal{H}^{\prime\prime}$ is a smooth section of $\widetilde{E}\boxtimes\widetilde{E}^{\prime}$ by the regularity of elliptic differential equations (see [hormander1983analysis]). ∎
@@
Remark 2.6.
In the framework of BCOV theory, $E=\wedge^{\bullet}T^{1,0}M$. The pairing $\omega$ is
$\omega(\alpha,\beta)=\operatorname{tr}(\alpha\wedge\beta),\qquad\alpha,\beta\in\Gamma(M,\wedge^{\bullet}T^{1,0}M),$
where $\operatorname{tr}$ denotes the trace map introduced in (1.1).
@@
The existence and uniqueness of the heat kernel follow from [berline2003heat]. The convergence $\lim_{t\to\infty}H_{t}=\mathcal{H}$ could be proved as in [berline2003heat, Section 2.6].
@@
The proof is similar to that of [berline2003heat, Theorem 2.38], and we outline the main steps.
@@
Since the singularities of the right-hand side are supported on the diagonal $\triangle$, it follows from elliptic regularity (cf. [hormander1983analysis]) that $\lim_{\epsilon\to 0}G_{\epsilon,L}$ is smooth off $\triangle$. Hence, so is
@@
By [berline2003heat, Proposition 2.37], the integral
@@
This follows directly from [berline2003heat, Proposition 2.37]. ∎
@@
Hence, by 2.10, the singularities of $P$ are fully determined by the singularities of $P_{t}$ at $t=0$.
@@ 2.2 Feynman graph integrals
2.12 is the most technically involved result in this paper. The key ingredients for its proof are presented in section 3 and section 4, while the complete proof is given in section 4.2.1. In this section, we assume the validity of 2.12 and use it to construct Feynman graph integrals.
@@
The analyticity assumption in 2.12 is essential for our proof. It remains an interesting question whether this theorem holds without it.
@@
For a coordinate-free and equivalent definition of holomorphic Lagrangian densities, see [Williams:2018ows, Definition 2.15].
@@
Let $(M,E,\omega)$ be a triple as in 2.12. Let $\vec{\Gamma}$ be a directed graph. Then
@@
Let $(M,E,\omega)$ be triples as in 2.12. Let $\vec{\Gamma}$ be a directed graph. The Feynman graph integral on $M$ is defined by:
@@
When $M$ is an elliptic curve and $E$ is the trivial vector bundle, the Feynman graph integrals can be computed explicitly, see [yang2024feynman].
In the remainder of this subsection, we introduce an alternative formulation of 2.21.
@@
Let $(M,E,\omega)$ be triples as in 2.12. Let $\vec{\Gamma}$ be a directed graph, and $\tilde{\rho}$ a fake distance function. Then we have